OpenStudy (anonymous):

What is the degree measure of an arc 4(pi) ft. long in a circle of radius 10 ft.?

OpenStudy (jh3power):

arc=2pi r times theta/360 theta=arc*360/2pi r theta=4pi/2pi *360/10 theta=72

OpenStudy (jh3power):

theta=degree measure